Can somebody nswer this question its easy well not 4 me i'm in 7th grade S=C+xC solve each equation for x
Answer by Edwin McCravy(20064)



      S = C + xC 

Add - C to both sides:


  S - C = C + xC - C

The C and the - C cancel out on the right

  S - C = xC

Divide both sides by C

 S - C     xC
------- = ----
   C        C

The C's cancel on the right:

 S - C     
------- = x
   C        

That's the answer.

Edwin